Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini - 15:31 Friday 12 April 2024 (63946) Print this report
PR_TX_SET investigation

Today I've noticed the presence of an offset on PR_TX (Fig.1). The "ball" generated by the 2-D istogram of PR_TX_INPUT and B4_12/112MHz_mag is skewed at a positive angle, suggesting that the maximum of the error signal's fluctuation should be around 0.006 mA. Both 12 and 112 MHz sidebands show agreement on this, which isn't the case for PR_TY.

Interestingly, this offset was not there a week ago (Fig.2).

I've tried to slowly change the setpoint of PR_TX to recover this offset. This was beneficial for the sidebands power, obviously, but the effect on the sensitivity is negligible (Fig.3) and it's invisible with respect to its normal fluctuation.

The offset has been introduced online, nothing in the automation has changed. Given the negligible effect on the interferometer, I see no harm in leaving it online for the rest of this lock and observe any possible difference over a longer period of time.

Images attached to this report
Comments to this report:
boldrini - 22:14 Friday 12 April 2024 (63953) Print this report

I've re-checked the offset on PR_TX.

Many hours have passed, but it seems that the offset I've put on PR_TX is still good (Fig.1). Some very light skewedness can be observed on the sidebands, indicating that this offset might have become larger overtime, but I would rather not risk disturbing the lock with such a weak indication. If the lock holds, I will keep monitoring this offset over the week-end.

Images attached to this comment
boldrini - 15:46 Monday 15 April 2024 (63969) Print this report

The offset on PR_TX is still there.

It's hard to tell exactly whithout trying, but it still seems that the optimal setpoint falls between 0.004 and 0.006 mA, as it did last Friday

Images attached to this comment
Search Help